The salaries of the president and Vice President of a certain country total $652,600 a year. If the president makes $247,400 more than the Vice President, find each of their salaries

Found 2 solutions by ewatrrr, ikleyn:
Answer by ewatrrr(24785)


Hi
(president makes $247,400 more )than the Vice President " x"
Question states:
 (x + $247,400) + x = $652,600
               2x = $405,200
 Solve          x = $202,600, VP salary
 President salary = $450,00

$202,600 + $450,000 = $652,600 checks.
Wish You the Best in your Studies.


Answer by ikleyn(52778)

.

    x + y = 652600

    x - y = 247400

--------------------------  Add the equations and see what will happen.


In this way, you will find x first;  then you will be in position to find y.


This method is called ELIMINATION method.




Alternatively, you may use one-equation setup


    x + (x-247400) = 652600 




Informal rule says:  the larger value is half sum of the given two values;

the smaller value is half of the difference of the two given values.

Solved, answered and explained.